how will this fraction will be simplified?
11
----
33
 
 

Grade:

41 Answers

Viraj C Bukitagar
9 years ago
11 and 33 is both divisible by 11 (11*1=11,11*3=33) therefore the simplified answer is 1/3
Ansh Agrawal
9 years ago
11/33 is divisible by 3. so 11 will become 1 when divided and 33 will become 3 when divided. so it will become 1/3
Sreeraj R Nair
9 years ago
Well it is quite simple.Just simplify both numerator and denominator by 11   thats is the answer is 1/3
11*1=11 and11*3=33
Yugabrat Gogoi
9 years ago
its quite simple its 1/3
as 11 x 1=11 and 11 x 3=33
Yugabrat Gogoi
9 years ago
its quite simple its 1/3
as 11 x 1=11 and 11 x 3=33
Aryan Vyas
9 years ago
To simplify 11/33,first find a common factor and then divide both the numerator and the denominator by it.In this case we take 11 as the common factor and after dividing 11 and 33 by 11, we ge 1/3.
Aryan Vyas
9 years ago
To simplify 11/33,first find a common factor and then divide both the numerator and the denominator by it.In this case we take 11 as the common factor and after dividing 11 and 33 by 11, we ge 1/3.
Aryan Vyas
9 years ago
To simplify 11/33,first find a common factor and then divide both the numerator and the denominator by it.In this case we take 11 as the common factor and after dividing 11 and 33 by 11, we ge 1/3.
 
Aryan Vyas
9 years ago
To simplify 11/33,first find a common factor and then divide both the numerator and the denominator by it.In this case we take 11 as the common factor and after dividing 11 and 33 by 11, we ge 1/3.
Aryan Vyas
9 years ago
To simplify 11/33,first find a common factor and then divide both the numerator and the denominator by it.In this case we take 11 as the common factor and after dividing 11 and 33 by 11, we ge 1/3.
Nikhil kumar
8 years ago
for finding simplified fraction we find common factor of numerator and denominator and then divide the numerator and denominator by the factor.
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
 
Nikhil kumar
8 years ago
for finding simplified fraction we find common factor of numerator and denominator and then divide the numerator and denominator by the factor.
 
Nikhil kumar
8 years ago
 
for finding simplified fraction we find common factor of numerator and denominator and then divide the numerator and denominator by the factor.
 
Nikhil kumar
8 years ago
 
for finding simplified fraction we find common factor of numerator and denominator and then divide the numerator and denominator by the factor.
 
Nikhil kumar
8 years ago
for finding simplified fraction we find common factor of numerator and denominator and then divide the numerator and denominator by the factor.
 
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
Nikhil kumar
8 years ago
In this case the factor is 11, so on dividing it by 11 and 33 you get 1/3
Prajwal Kavad
8 years ago
answer is 1/3
Adarsh
8 years ago
The simplest form of 11/13 is 1/3
Ritika Das
6 years ago
The greatest common divisor of 11 and 33 is 11 by finding out the HCF of the two numbers.
Therefore we can cancel out 11/11=1 and consequently 33/11=3.
So the final simplified result will be 1/3. 
Harshit Panwar
6 years ago
11/33=1/3
Hope it helps you
C.Vishaal
6 years ago
11/33 If divided by 11(Both the numerator and the denominator) gives you an answer of 1/3.
 
Hope this helps.
Vivek
6 years ago
You can solve this by simply dividing both the numerator and denominator by he common factor of the bothe no. And that no. Is 11. So divide by 11. We gat 1/3
Priyansh
6 years ago
1/3 or 0.33333333333333333333333333333333333333333333333333333333333333333333333333333333333333333333333
Ishita
5 years ago
First of all divide the numerator and denominator of11/33 by 11 then you will get 1/3 which is the most simplified number means it can't be simplified more . Yes,you can convert it to decimal and answer will be 0.333....
 
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you 
 
 
Om
5 years ago
 
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you!!
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you!
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you :)
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you !!!!
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you :) :)
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you
 
with regards
your fellow student
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you!!!!!!!!
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you!!!!!!!!!!
Om
5 years ago
Dear Student
To simplify any fraction first we need to observe the possibilities for cancelling them
As per the given fraction 
 11   can be written as follows
 33
 
 11  =    11 
 33        33 
       =  1        (since 11 and 33 are divisible by 11)
           3   
I hope this answer helps you!!
 
with regards
your fellow student